Question 4 2 pts

An analysis of variances produces dftotal = 32 and dfwithin = 29. For this analysis, how many treatment conditions are being compared?

2

4

3
Answer #1

We can see that

Df total =32

Df within =29

Df total = df treatment + df within

Df treatment =32-29  

Df treatment =3

Now k = number treatment

Df treatment = k-1

K = df treatment +1

K = 3 + 1

K= 4

So number of treatment = 4